• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Sirca

    Log entry time 10:30:34 on October28,2001

    Entry number 75303

    keyword=Cryo problems due to communication failures solved

    The network problems affecting the ROCs earlier this morning
    seemed to influence also the IOCs, in particular the iocha26 which
    appears to be the IOC which controls the cryogen levels for the
    spectrometers. This in turn caused the slow controls to fail to
    control the flow of helium. In consequence, the magnets
    (Left D1 and Right Q3) tripped for lack of helium. Although the
    control screens in the counting room showed sufficient cryogen
    levels, they did not update. The magnet power supplies were then
    reset in the hall and the IOC 26 was rebooted, and we could recover
    the magnets. But we could not restore communication to iocha26,
    although we pushed both E_COM_RST and H_COM_RST.
    The screen with the IOC heartbeats shows an empty slot for
    iocha26 although MCC tells us it is alive.

    We are now restoring magnet currents to their required values
    for physics production.